The length of a rectangular yard is 10 feet longer than half its width. The perimeter is 47 ft. What is the area?
Answer by checkley77(12844)

L=10+W/2
P=2L+2W
47=2(10+W/2)+2W
47=20+2W/2+2W
47-20=W+2W
27=3W
W=27/3
W=9 ANS FOR THE WIDTH.
L=10+9/2
L=10+4.5=14.5 ANS. FOR THE LENGTH.
A=L*W
A=14.5*9
130.5 ANS FOR THE AREA.